OpenStudy (anonymous):

What is the slope (m) and the y-intercept (b) of the line y+4=0? A) m=-4, b=0 B) m=0, b=-4 C) m=1, b=-4 D) m=1, b=4

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

get y alone first :)

OpenStudy (anonymous):

y=mx+b form?

jigglypuff314 (jigglypuff314):

yeah, just isolate y subtract both sides by 4

jigglypuff314 (jigglypuff314):

y+4=0 subtract 4 from both sides -4 -4 y = -4 you can look at it like y = 0x - 4

OpenStudy (anonymous):

so, its A?

jigglypuff314 (jigglypuff314):

not quite :)

OpenStudy (anonymous):

:(

jigglypuff314 (jigglypuff314):

y = mx + b y = 0x - 4

OpenStudy (anonymous):

so, I add 4?

jigglypuff314 (jigglypuff314):

no, you just see how they line up and decide what m and b equal

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m=0 b=4

OpenStudy (anonymous):

B?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Thank you!

jigglypuff314 (jigglypuff314):

you're welcome :)